OpenStudy (anonymous):

If the ratio of two supplementary angles is 7 to 8, calculate the measure of the larger of the two

OpenStudy (lgbasallote):

the ratio is 7:8 and the total is 180 7x + 8x = 180 15x = 180 x = 180/15 x = 12 so obviously 8x is the larger angle so sub x into that 8x = 8(12) got it @JeremiahJallo

OpenStudy (anonymous):

Thank you @lgbasallote

OpenStudy (lgbasallote):

welcome